A sample of 10 networking sites for a specific month has a mean of 26.1 and a standard deviation of 4.2. Find the 99% confidence interval of the true mean.


1
Expert's answer
2021-07-14T18:14:39-0400

"n=10 \\
\bar{x} =26.1 \\
s=4.2"

Two-sided confidence interval:

"CI = (\bar{x} - \frac{Z_c \times s}{\sqrt{n}}, \bar{x} + \frac{Z_c \times s}{\sqrt{n}}) \\

CI = (26.1 - \frac{2.576 \times 4.2}{\sqrt{10}}, 26.1 + \frac{2.576 \times 4.2}{\sqrt{10}}) \\

CI = (26.1 - 3.42, 26.1 + 3.42) \\

CI = (22.68, 29.52)"